# Plain Transformers Can be Powerful Graph Learners

### v2

@@ D.2 MPNNs
As the most widely used GNNs, message-passing neural networks (MPNNs) Gilmer et al., (2017); Kipf and Welling, (2017); Hamilton et al., (2017); Veličković et al., (2018) learn graphs, following the 1-WL framework Xu et al., (2019). However, there are several known limitations of MPNNs: (1) over-smoothing Li et al., (2018); Oono and Suzuki, (2020); (2) over-squashing and under-reaching Alon and Yahav, (2020); Topping et al., (2022); and (3) limited expressivity bounded by $1$-WL Xu et al., (2019).
@@ D.3 Graph Positional Encoding and Structural Encoding
Attention mechanisms are structure-invariant operators which sense no structural information inherently. Therefore, Transformers strongly rely on positional encoding to capture the structural information Vaswani et al., (2017); Dosovitskiy et al., (2021); Su et al., (2024). Designing graph positional encoding is challenging compared to the counterpart in Euclidean spaces, due to the irregular structure and the symmetry to permutation. Widely used graph positional encoding includes absolute PE: LapPE Dwivedi and Bresson, (2021); Huang et al., (2024); and relative PE: shortest-path distance Ying et al., (2021), resistance-distance Zhang et al., 2023b , and RRWP Ma et al., (2023). Recent works Black et al., (2024); Zhang et al., 2024b study the connections between absolute PE and relative PE.
Besides the aforementioned PE, there exist several structural encoding (SE) approaches that aim to enhance MPNNs, e.g., RWSE Dwivedi et al., 2022b , substructure counting Bouritsas et al., (2022), and homomorphism counting Jin et al., (2024). These SEs can effectively improve the empirical performance and/or theoretical expressivity of MPNNs. Although they are not designed specifically for graph Transformers, integrating them into graph Transformers is usually beneficial.
@@ D.4 Higher-order GNNs
Besides graph Transformers, inspired by the $K$-WL Weisfeiler and Leman, (1968) and $K$-Folklore WL Cai et al., (1992) frameworks, $K$-GNNs Morris et al., (2019); Zhou et al., (2023); Maron et al., (2019) uplift GNNs to higher-order, treating a tuple of $K$ nodes as a token and adapting the color refinement algorithms accordingly. Some other higher-order GNNs Bodnar et al., (2022, 2021) are less closely related to $K$-WL but still perform well. $K$-GNNs can reach theoretical expressivity bounded by $K$-WL, but are typically computationally costly, with $O(N^{K})$ computational complexity.
@@ D.5 Subgraph GNNs
Subgraph GNNs Bevilacqua et al., (2022); Frasca et al., (2022); Zhang et al., 2024a ; Zhang and Li, (2021); Huang et al., (2023) focus on improving the expressivity of GNNs beyond 1-WL. Unlike graph Transformers and higher-order GNNs, subgraph GNNs typically do not change the model architecture. Instead, MPNNs are trained using a novel learning pipeline: a graph is split into multiple subgraphs, subgraph representations are learned independently, and then the subgraph representations are merged to obtain the graph representation. Node-level subgraph GNNs are typically bounded by $3$-WL Frasca et al., (2022), while the expressivity of edge-level subgraph GNNs is still under-explored. Huang et al., (2023) reveal that the expressivity of edge-level GNNs might partially go beyond $3$-WL. Note that subgraph GNNs usually have high memory requirements due to the need to save multiple copies for each input graph.
@@ E.1 Theoretical Expressivity of PPGT
The detailed discussion of GD-WL GNNs and its theoretical expressivity can be found in Sec. 4 and Appx. E.3 of Zhang et al., 2023b , as well as in Sec. 5 of Zhang et al., 2024b . The discussion on RRWP can be found in Sec. 3.1.1 of Ma et al., (2023). The empirical study of realized expressivity (in Sec. 5.1) also matches the theoretical expressivity of PPGT.
Based on the GD-WL analysis framework Zhang et al., 2023b , the proof is straightforward. However, we still provide a simple proof for the completeness of the conclusion.
@@
(Lemma 5 of  Xu et al., (2019)) For any countable set $\mathcal{X}$, there exists a function $f:\mathcal{X}\to\mathbb{R}^{n}$ such that $h(\hat{\mathcal{X}}):=\sum_{x\in\hat{\mathcal{X}}}f(x)$ is unique for each multiset $\hat{\mathcal{X}}\in\mathcal{X}$ of bounded size. Moreover, for some function $\phi$, any multiset function $g$ can be decomposed as $g(\hat{\mathcal{X}})=\phi(\sum_{x\in\hat{\mathcal{X}}}f(x))$.
@@
In this proof, we consider shortest-path distance (SPD) as an example of generalized distance (GD), denoted as $\text{d}_{G}^{\text{SPD}}$, which can be directly extended to other GDs such as the resistance distance (RD) Zhang et al., 2023b and RRWP Ma et al., (2023). Note that the choice of GD determines the practical expressiveness of GD-WL.
@@
Then the GD-WL aggregation at the $\ell$-th iteration in Eq. (13) can be equivalently rewritten as (See Theorem E.3 in Zhang et al., 2023b ):
@@
Assuming the color of each node $\chi^{l}_{G}(v)$ is represented as a vector $\mathbf{x}_{v}^{(l)}\in\mathbb{R}^{C}$, by setting the query and key projection matrices $\mathbf{W}_{Q},\mathbf{W}_{K}$ as zero matrices and $\theta$ as a zero function following Zhang et al., 2023b , the attention layer of PPGT with URPE of the $h$-th head (Eq. 5) can be written as
@@
It is worth mentioning that the expressivity upper bound of GD-WL (i.e., 2-FWL/3-WL), given by Theorem 4.5 in Zhang et al., 2023b , is based on the usage of generalized distance (e.g., resistance distance) as graph PE. If we add PE that provides finer structural information, PPGT might surpass the 3-WL expressivity upper bound. This is empirically verified in Section 5.1, where we demonstrate using I2GNN to generate PE for PPGT.
@@ E.2 LN and RMSN are Magnitude Invariant
$\text{RMSN}(c\cdot\mathbf{x})$ $:=\frac{c\cdot\mathbf{x}}{\frac{1}{\sqrt{D}}\|c\cdot\mathbf{x}\|}\cdot\boldsymbol{\gamma}\,,$
@@ Appendix F Limitations and Discussion
As a plain Transformers, PPGTs still require $O(N^{2})$ computational complexity like other plain Transformers.
@@
  • Large-scale-graph dataset (many nodes in a graph): This corresponds to the challenges in long-context learning in language tasks and high-resolution image processing in vision tasks. The challenges mainly lie in the computational efficiency and memory consumption of the models. These datasets are usually small-scale in terms of examples, typically containing only one graph per dataset.
@@ F.2 Efficiency Techniques for Plain Transformers
BigBird Zaheer et al., (2020) and Exphormer Shirzad et al., (2023) propose building sparse attention mechanisms through sparse computational graphs, using random graphs and expander graphs for language tasks and graph learning tasks, respectively. They can be directly adopted by PPGTs.
@@
Ainslie et al., (2023) propose grouped-query attention (GQA) to save the memory bandwidth. It is a more specific design for decoder-only Transformers (for KV-cache), and might not introduce remarkable benefits to encoder-based Transformers, such as ViTs Dosovitskiy et al., (2021); Touvron et al., 2021a ; Liu et al., (2021), Time-series Transformers Nie et al., (2023); Zhang et al., 2024c and most GTs Kreuzer et al., (2021); Ying et al., (2021); Zhang et al., 2023b ; Ma et al., (2023) including PPGTs.
@@ F.3 Impact Statement
This paper presents work whose goal is to advance the field of Geometric/Graph Deep Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.

### v3

@@ D.2 MPNNs
As the most widely used GNNs, message-passing neural networks (MPNNs) Gilmer et al. (2017); Kipf and Welling (2017); Hamilton et al. (2017); Veličković et al. (2018) learn graphs, following the 1-WL framework Xu et al. (2019). However, there are several known limitations of MPNNs: (1) over-smoothing Li et al. (2018); Oono and Suzuki (2020); (2) over-squashing and under-reaching Alon and Yahav (2020); Topping et al. (2022); and (3) limited expressivity bounded by $1$-WL Xu et al. (2019).
@@ D.3 Graph Positional Encoding and Structural Encoding
Attention mechanisms are structure-invariant operators which sense no structural information inherently. Therefore, Transformers strongly rely on positional encoding to capture the structural information Vaswani et al. (2017); Dosovitskiy et al. (2021); Su et al. (2024). Designing graph positional encoding is challenging compared to the counterpart in Euclidean spaces, due to the irregular structure and the symmetry to permutation. Widely used graph positional encoding includes absolute PE: LapPE Dwivedi and Bresson (2021); Huang et al. (2024); and relative PE: shortest-path distance Ying et al. (2021), resistance-distance Zhang et al. (2023b), and RRWP Ma et al. (2023). Recent works Black et al. (2024); Zhang et al. (2024b) study the connections between absolute PE and relative PE.
Besides the aforementioned PE, there exist several structural encoding (SE) approaches that aim to enhance MPNNs, e.g., RWSE Dwivedi et al. (2022b), substructure counting Bouritsas et al. (2022), and homomorphism counting Jin et al. (2024). These SEs can effectively improve the empirical performance and/or theoretical expressivity of MPNNs. Although they are not designed specifically for graph Transformers, integrating them into graph Transformers is usually beneficial.
@@ D.4 Higher-order GNNs
Besides graph Transformers, inspired by the $K$-WL Weisfeiler and Leman (1968) and $K$-Folklore WL Cai et al. (1992) frameworks, $K$-GNNs Morris et al. (2019); Zhou et al. (2023); Maron et al. (2019) uplift GNNs to higher-order, treating a tuple of $K$ nodes as a token and adapting the color refinement algorithms accordingly. Some other higher-order GNNs Bodnar et al. (2022; 2021) are less closely related to $K$-WL but still perform well. $K$-GNNs can reach theoretical expressivity bounded by $K$-WL, but are typically computationally costly, with $O(N^{K})$ computational complexity.
@@ D.5 Subgraph GNNs
Subgraph GNNs Bevilacqua et al. (2022); Frasca et al. (2022); Zhang et al. (2024a); Zhang and Li (2021); Huang et al. (2023) focus on improving the expressivity of GNNs beyond 1-WL. Unlike graph Transformers and higher-order GNNs, subgraph GNNs typically do not change the model architecture. Instead, MPNNs are trained using a novel learning pipeline: a graph is split into multiple subgraphs, subgraph representations are learned independently, and then the subgraph representations are merged to obtain the graph representation. Node-level subgraph GNNs are typically bounded by $3$-WL Frasca et al. (2022), while the expressivity of edge-level subgraph GNNs is still under-explored. Huang et al. (2023) reveals that the expressivity of edge-level GNNs might partially go beyond $3$-WL. Note that subgraph GNNs usually have high memory requirements due to the need to save multiple copies for each input graph.
@@ E.1 Theoretical Expressivity of PPGT
The detailed discussion of GD-WL GNNs and its theoretical expressivity can be found in Sec. 4 and Appx. E.3 of Zhang et al. (2023b), as well as in Sec. 5 of Zhang et al. (2024b). The discussion on RRWP can be found in Sec. 3.1.1 of Ma et al. (2023). The empirical study of realized expressivity (in Sec. 5.1) also matches the theoretical expressivity of PPGT.
Based on the GD-WL analysis framework Zhang et al. (2023b), the proof is straightforward. However, we still provide a simple proof for the completeness of the conclusion.
@@
(Lemma 5 of  Xu et al. (2019)) For any countable set $\mathcal{X}$, there exists a function $f:\mathcal{X}\to\mathbb{R}^{n}$ such that $h(\hat{\mathcal{X}}):=\sum_{x\in\hat{\mathcal{X}}}f(x)$ is unique for each multiset $\hat{\mathcal{X}}\in\mathcal{X}$ of bounded size. Moreover, for some function $\phi$, any multiset function $g$ can be decomposed as $g(\hat{\mathcal{X}})=\phi(\sum_{x\in\hat{\mathcal{X}}}f(x))$.
@@
In this proof, we consider shortest-path distance (SPD) as an example of generalized distance (GD), denoted as $\text{d}_{G}^{\text{SPD}}$, which can be directly extended to other GDs such as the resistance distance (RD) Zhang et al. (2023b) and RRWP Ma et al. (2023). Note that the choice of GD determines the practical expressiveness of GD-WL.
@@
Then the GD-WL aggregation at the $\ell$-th iteration in Eq. (13) can be equivalently rewritten as (See Theorem E.3 in Zhang et al. (2023b)):
@@
Assuming the color of each node $\chi^{l}_{G}(v)$ is represented as a vector $\mathbf{x}_{v}^{(l)}\in\mathbb{R}^{C}$, by setting the query and key projection matrices $\mathbf{W}_{Q},\mathbf{W}_{K}$ as zero matrices and $\theta$ as a zero function following Zhang et al. (2023b), the attention layer of PPGT with URPE of the $h$-th head (Eq. (5)) can be written as
@@
It is worth mentioning that the expressivity upper bound of GD-WL (i.e., 2-FWL/3-WL), given by Theorem 4.5 in Zhang et al. (2023b), is based on the usage of generalized distance (e.g., resistance distance) as graph PE. If we add PE that provides finer structural information, PPGT might surpass the 3-WL expressivity upper bound. This is empirically verified in Section 5.1, where we demonstrate using I2GNN to generate PE for PPGT.
@@ E.2 LN and RMSN are Magnitude Invariant
$\text{RMSN}(c\cdot\mathbf{x})$ $:=\frac{c\cdot\mathbf{x}}{\frac{1}{\sqrt{D}}\|c\cdot\mathbf{x}\|}\cdot\bm{\gamma}\,$,
@@ Appendix F Limitations and Discussion
As plain Transformers, PPGTs still require $O(N^{2})$ computational complexity like other plain Transformers.
@@
  • Large-scale-graph dataset (many nodes in a graph): This corresponds to the challenges in long-context learning in language tasks and gigapixel image processing in vision tasks. The challenges mainly lie in the computational efficiency and memory consumption of the models. These datasets are usually small-scale in terms of examples, typically containing only one graph per dataset.
@@ F.2 Efficiency Techniques for Plain Transformers
BigBird Zaheer et al. (2020) and Exphormer Shirzad et al. (2023) propose building sparse attention mechanisms through sparse computational graphs, using random graphs and expander graphs for language tasks and graph learning tasks, respectively. They can be directly adopted by PPGTs.
@@
Ainslie et al. (2023) proposes grouped-query attention (GQA) to save the memory bandwidth. It is a more specific design for decoder-only Transformers (for KV-cache), and might not introduce remarkable benefits to encoder-based Transformers, such as ViTs Dosovitskiy et al. (2021); Touvron et al. (2021a); Liu et al. (2021), Time-series Transformers Nie et al. (2023); Zhang et al. (2024c) and most GTs Kreuzer et al. (2021); Ying et al. (2021); Zhang et al. (2023b); Ma et al. (2023) including PPGTs.
@@ F.3 Impact Statement
This paper presents work whose goal is to advance the field of Geometric/Graph Deep Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.